# Contrastive Forward-Forward: A Training Algorithm of Vision Transformer

### v1

@@ 4.4 How CFF Stacks Up Against Backpropagation
Random Crop CIFAR10 ViT[128 4 5] 76.92 80.42
@@
Random Crop + RandAug CIFAR10 ViT[128 4 5] 86.33 81.73
@@
(a) Augmention: Random Crop
(b) Augmention: RandAug [45]
@@
Further analysis of Table 3 reveals a specific pattern in comparing BP+CE with the proposed algorithm concerning model complexity and the type of augmentation used. It seems that CFF+M demonstrates higher robustness to overfitting compared to BP as model complexity increases. To illustrate this, Fig. 7 shows accuracy vs. model complexity under two augmentation scenarios. First, it is observed that using the proposed loss in the method (CFF+M) leads to better performance compared to using the original loss in the method (CFF) across all cases. Analysis of these plots shows that CFF+M tends to overfit at higher complexities compared to BP. This behavior can be attributed to an inherent form of regularization in FF-based method which layers have light collaboration. It can be particularly useful in applications where there is a higher risk of overfitting, such as when powerful augmentations like RandAug [45] are not employed or lack of training data. Furthermore, when using RandAug, which serves as the primary playing field for comparison, the proposed method can achieve competitive performance at higher complexities, which was obtained with BP in lower model complexity scenario.
@@
50% 0% 78.27 (8.68$\downarrow$)↓ ) 80.58 (5.59$\downarrow$)↓ )
80% 0% 69.19 (17.76$\downarrow$)↓ ) 73.26 (12.91$\downarrow$)↓ )
0% 20% 78.10 (8.85$\downarrow$)↓ ) 78.33 (6.84$\downarrow$)↓ )
0% 40% 70.36 (16.59$\downarrow$)↓ ) 71.20 (13.97$\downarrow$)↓ )
@@ 5.1 Layer Training Parallelization
Fig. 8 illustrates the proposed pipeline. In this figure, an example with a 4-layer network using 4 GPUs is considered. For simplicity, the one-forward approach is used (Sec. 4.5). The entire figure pertains to one epoch, where each step is an iteration along the batches. At $t=1$, the first batch enters the first layer. The local forward and local backward processes in this layer are performed by GPU 1. The output of this layer is transferred to the second layer at $t=2$. Since the first layer and GPU 1 are idle, they can take in the second batch, while at the same time, GPU 2 processes the second layer. As a result, in addition to preparing the output of the second batch in the first layer, the output of the second layer belonging to the first batch is also ready; And following the same procedure up to $t=4$. At $t=4$, the layers are processing representatives from layers 1 to 4 for batches 4 to 1, respectively and there are effectively 4 batches within the network. In this case after $t\geq 4$, the output for a batch is ready at each step from the viewpoint of the whole network. Therefore, if the total number of batches is $N$ and the total number of required steps is $T$, the first batch requires 4 steps to pass through the entire network, while the output for the second batch is ready at $t=5$, and similarly, the output for subsequent batches is ready at each step. Hence, $T=4+(N-1)$. For BP, each batch requires 4 steps, so $T=4N$, which means that this parallelization, approximately speeds up training by $4$ in each epoch. Fortunately, each transformer layer shown in Fig. 3 has the same structure. Consequently, it is expected that processing time at each layer with a fixed batch size will be equal, which reinforces the assumptions made. However, a challenge that is not considered in this example is the IO processing time. Since batches in typical training methods are read sequentially from storage, this can create a processing time bottleneck.
@@ 6 Conclusion
FF [3] is a new method for training neural networks, specifically designed for image classification. This algorithm calculates a loss function after each layer, making the training of the layers independent of each other, except for preparing their inputs. This approach is different from the traditional BP algorithm, which employs the CE loss function (but is not limited to this loss) at the end of the network. From this point of view, FF is more similar to brain function, where it has been shown that BP is not [1].
FF uses two forward passes: one with the correct input label and another with an incorrect label. Layers are trained by comparing the two resulting representations. This comparison between the two representations can also be found in another well-known field named Contrastive Learning [15, 16, 17, 19]. Therefore, we used insights from it to revise FF. We replaced the original FF loss function with a supervised contrastive loss and adapted the input sample feeding method to align with SCL [19]. Thus, the proposed method can be viewed from two points of view: as SCL where the loss function is placed after each layer instead of at the end (similar to FF), or as FF with contrastive loss.
@@
The main advantage of using FF-based methods (including the proposed method) is the ability to create a pipeline that allows strong parallelization during training across layers. The proposed algorithm (inherited from FF) has some disadvantages compared to BP including the limited choice of loss functions, relatively complex implementation, and restriction to classification problem. We hope that researchers will pay more attention to this algorithm and address its shortcomings.

### v2

@@ 4.4 How CFF Stacks Up Against Backpropagation
Random Crop
CIFAR10 ViT[128 4 5] 76.92 80.42
@@
Random Crop + RandAug
CIFAR10 ViT[128 4 5] 86.33 81.73
@@
(a) Augmentation: Random Crop
(b) Augmentation: RandAug [45]
@@
Further analysis of Table 3 reveals a specific pattern in comparing BP+CE with the proposed algorithm concerning model complexity and the type of augmentation used. It seems that CFF+M demonstrates higher robustness to overfitting compared to BP as model complexity increases. To illustrate this, Fig. 7 shows accuracy vs. model complexity under two augmentation scenarios. First, it is observed that using the proposed loss in the method (CFF+M) leads to better performance compared to using the original loss in the method (CFF) across all cases. Analysis of these plots shows that CFF+M tends to overfit at higher complexities compared to BP. This behavior can be attributed to an inherent form of regularization in FF-based method which layers have light collaboration. It can be particularly useful in applications where there is a higher risk of overfitting, such as when powerful augmentations like RandAug [45] are not employed or lack of training data. Furthermore, when using RandAug, which serves as the primary playing field for comparison, the proposed method can achieve competitive performance at higher complexities, which was obtained with BP in a lower model complexity scenario.
@@
[…]
@@ 5.1 Layer Training Parallelization
Fig. 8 illustrates the proposed pipeline. In this figure, an example with a 4-layer network using 4 GPUs is considered. For simplicity, the one-forward approach is used (Sec. 4.5). The entire figure pertains to one epoch, where each step is an iteration along the batches. At $t=1$, the first batch enters the first layer. The local forward and local backward processes in this layer are performed by GPU 1. The output of this layer is transferred to the second layer at $t=2$. Since the first layer and GPU 1 are idle, they can take in the second batch, while at the same time, GPU 2 processes the second layer. As a result, in addition to preparing the output of the second batch in the first layer, the output of the second layer belonging to the first batch is also ready; And following the same procedure up to $t=4$. At $t=4$, the layers are processing representatives from layers 1 to 4 for batches 4 to 1, respectively and there are effectively 4 batches within the network. In this case after $t\geq 4$, the output for a batch is ready at each step from the viewpoint of the whole network. Therefore, if the total number of batches is $N$ and the total number of required steps is $T$, the first batch requires 4 steps to pass through the entire network, while the output for the second batch is ready at $t=5$, and similarly, the output for subsequent batches is ready at each step. Hence, T=4+(N1)T=4+(N-1). For BP, each batch requires 4 steps, so $T=4N$, which means that this parallelization, approximately speeds up training by $4$ in each epoch. Fortunately, each transformer layer shown in Fig. 3 has the same structure. Consequently, it is expected that processing time at each layer with a fixed batch size will be equal, which reinforces the assumptions made. However, a challenge that is not considered in this example is the IO processing time. Since batches in typical training methods are read sequentially from storage, this can create a processing time bottleneck.
@@ 6 Conclusion
FF [3] is a novel method for training neural networks, specifically designed for image classification. This algorithm calculates a loss function after each layer, making the training of the layers independent of each other, except for preparing their inputs. This approach is different from the traditional BP algorithm, which employs the CE loss function (but is not limited to this loss) at the end of the network. From this point of view, FF is more plausible to brain function, where it has been shown that BP is not [1].
FF uses two forward passes: one with the correct input label and another with an incorrect label. Layers are trained by comparing the two resulting representations. This comparison between the two representations can also be found in another well-known field named Contrastive Learning [15, 16, 17, 19]. Therefore, we used insights from it to modify FF. We replaced the original FF loss function with a supervised contrastive loss and adapted the input sample feeding method to align with SCL [19]. Thus, the proposed method can be viewed from two points of view: as SCL where the loss function is placed after each layer instead of at the end (similar to FF), or as FF with contrastive loss.
@@
The proposed algorithm has some disadvantages compared to BP including the limited choice of loss functions, relatively complex implementation, and restriction to classification problem. Extending the method to complex vision tasks like object tracking/detection and image segmentation, evaluating it on large-scale models and bigger datasets, and examining its biological plausibility are suggested as future work.